# The Mollified (Discrete) Uniform Distribution and its Applications

### v1

@@ 5 Conclusions
In this article, the (continuous) mollified uniform distribution was rediscovered, and a novel discrete counterpart was proposed as well. Both types of mollified uniform distribution are defined as the convolution of the (discrete) uniform distribution and a centered symmetric mollifier. Relevant stochastic properties are computed by elementary calculations such that the mollified uniform distributions might serve as neat examples in teaching. The mollified uniform distribution also constitutes a model that may cover a wide range of kurtosis levels. Furthermore, its cdf and qf, respectively, are used for the defining soft-clipping regression models, which can be adapted to binary, ordinal, and bounded quantitative outcomes. Here, a couple of future research directions become clear. Up to now, the soft-clipping GLM approach was only investigated in rather special setups, namely as the output function of an ANN in order to generate bounded real-valued outcomes in the unit interval $(0;1)$, see Klimek & Perelstein (2020), or for modeling time series of bounded counts (see Weiß & Jahn, 2022). But applications to, e. g., ordinary count regression (without the time aspect), to binary data, or ordinal data (as cumulative GLMs) appear to not have been considered so far.

### v2

@@ 5 Conclusions
In this article, the (continuous) mollified uniform distribution was rediscovered, and its origin as well as possible applications were reviewed. In addition, a novel discrete mollified uniform distribution was proposed, which uses an analogous construction as its continuous counterpart, namely a convolution of the (discrete) uniform distribution and a centered symmetric mollifier. Relevant stochastic properties can be computed by elementary calculations such that the mollified uniform distributions might also serve as neat examples in teaching. From a probabilistic point of view, the mollified uniform distribution constitutes a model that may cover a wide range of kurtosis levels. Furthermore, its cdf and qf, respectively, are used for the defining soft-clipping regression models, which can be adapted to binary, ordinal, and bounded quantitative data and time series. Here, the main feature of the soft-clipping models is given by their ability (provided that the scale parameter $c$ is sufficiently close to zero) to nearly behave like linear models (and thus being easy to interpret) while avoiding restrictive parameter constraints and being able to handle negative dependencies.
Let us conclude by outlining a couple of future research perspectives. Up to now, the soft-clipping GLM approach was only investigated in rather special setups, namely as the output function of an ANN in order to generate bounded real-valued outcomes in the unit interval $(0;1)$, see Klimek & Perelstein (2020), or for modeling bounded discrete-valued time series (see Weiß & Jahn, 2024). But applications to ordinary discrete-valued regression (without the time aspect) have not been considered so far, so soft-clipping GLMs for binary data, ordinal data (as cumulative GLMs), and count data appear to be important directions for future research. In particular, soft-clipping GLMs for bounded continuously distributed phenomena deserve attention both for traditional and time-series setups. Finally, the mollified uniform distribution could be an interesting choice for Bayesian statistics, where, similar to Yang & Rannala (2006), it could serve as a kind of uninformative prior without hard bounds.
Conflicts of Interest
The author declares no conflicts of interest.
Data Availability Statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.